the sum of the roots of a quadratic equation is -4. if one of the roots is 7, how would you determine the equation? write the equation.
Answer by Edwin McCravy(20054)

By this formula:

 

Put -4 for the sum of roots.



One root is given as 7.
Let the other root be R:

Then 7 + R = -4
         R = -4 - 7
         R = -11

Put (7)(-11)=-77 for the product of roots:

 

Simplify the signs:

 

-------------------------

Check by factoring:

     x² + 4x - 77 = 0
  (x + 11)(x - 7) = 0

x + 11 = 0;   x - 7 = 0
     x = -11;     x = 7     

Edwin